Question 1:

A nuclear power plant operates at 70 % of its maximum theoretical (Carnot) efficiency between temperatures of 660 ∘C and 340 ∘C. If the plant produces electric energy at the rate of 1.3 GW , how much exhaust heat is discharged per hour?

Question 2:

How long does it take the Sun to melt a block of ice at 0∘C with a flat horizontal area 1.0 m2 and thickness 2.0 cm ? Assume that the Sun's rays make an angle of 35 ∘ with the vertical and that the emissivity of ice is 0.050.

Question 3:

How much power is radiated by a tungsten sphere (emissivity ϵ = 0.35) of radius 17 cm at a temperature of 19 ∘C?

If the sphere is enclosed in a room whose walls are kept at -5 ∘C, what is the net flow rate of energy out of the sphere?
Answer #1

ONE QUESTION AT A TIME PLEASE............DO NOT POST MORE THAN ONE QUESTION AT A TIME

Question 1

Efficiency = 933 - 613 / 933 = 0.3429 ( this is theoretical efficiency)

actual efficiency = 0.7 * 0.3429 = 0.24

total power = 1.3 / 0.24 = 5.416 W

so,

power wasted = 4.1166 W

W = 4.1166 * 3600 = 14820 GJ/hr

or

W = 1.48e13

----------------------------------------------------------------------------------------

Question 2

Heat released

Q = ice * A * t *Lf

so,

time T = ice * A * t *Lf / e I A cos 35

WHERE e is efficiency, I is solar energy input (1000 W/m2)

T = 917 * 0.02 * 3.33e5 / 0.05*1000 * cos 35

T = 149110 seconds

-------------------------------------------------------------------------------